Simplify and express each of the following in exponential form:
20 + 30 + 40

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

= 1 + 1+ 1
= 3

Thanks
Shraddha Singh
2 years ago
Dear Student,
We know that any number raised to power zero is 1.
 
So, 2+ 3+ 4 = 1 + 1 + 1 = 3
 
Hope it helps!
Thank you ✌️